OpenStudy (anonymous):

1) What is the range of f(x) = -x2 + 4, if the domain is {2,0,1}? A) {0,3,5} B) {0,4,3} C) {0,4,5} D) {8,4,3} @hero @thesmartone @michele_lanio

OpenStudy (michele_laino):

the requested range is given by the subsequent subset of the real line: \[\huge range = \left\{ {f\left( 0 \right),\quad f\left( 1 \right),\quad f\left( 2 \right)} \right\}\]

OpenStudy (michele_laino):

so, we have to compute these quantities: \[\huge \begin{gathered} f\left( 0 \right) = - {\left( 0 \right)^2} + 4 = ...? \hfill \\ \hfill \\ f\left( 1 \right) = - {\left( 1 \right)^2} + 4 = ...? \hfill \\ \hfill \\ f\left( 2 \right) = - {\left( 2 \right)^2} + 4 = ...? \hfill \\ \end{gathered} \]

OpenStudy (anonymous):

sooo

OpenStudy (anonymous):

im still confused

OpenStudy (michele_laino):

what is: \(0 \cdot 0+4=...?\)

OpenStudy (michele_laino):

we have \(0 \cdot 0 +4=4\) so \(f(0)=-(0\cdot 0 )+4=4\) then \(y=4\) belong to the requested range now, please do the same with \(f(1)\)

OpenStudy (michele_laino):

belongs*

OpenStudy (anonymous):

so 0,4,5

OpenStudy (michele_laino):

\(y=0,\;y=4\) are correct! whereas we have: \(f(1)=-(1 \cdot 1)+4=-1+4=...?\)

OpenStudy (anonymous):

0,3

OpenStudy (michele_laino):

it si \(f(1)=3\), so the requested range is the set \(y=0\) \(y=3\) and \(y=4\) so, what is the right option?

OpenStudy (anonymous):

b

OpenStudy (michele_laino):

that's right!

OpenStudy (anonymous):

cant u help me on the next ones

OpenStudy (michele_laino):

ok!